a soap bubble appears green at the point on its front surface nearest the viewer. what is the smallest thichness of the sapbubble film could have? N=1.35

Grade:12

1 Answers

raghav sharma
12 years ago

according to interfernce:
if soap bubble appears green then green light should have maxima at that point
and as the light is returning to the viewer so it attains additional path difference of 180degree
so the condition becomes:

dX=(2n-1)L/2 INSTEAD OF dX=(n)L where L=wavelenght of green light

and dX= 2ut where u=refractive index of soap and t=thickness of soap film

so,2ut=(2n-1)L/2

as we have to find minimum distance so,

t=L/4u keeping n=1

keep u=1.35 L=WAVELENGTH OF GREEN LIGHT

HOPE IT IS CORRECT